OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (callisto):

When you have two points, you can first find the slope m by \[m = \frac{y_1-y_0}{x_1-x_0}\] Then, you can form the equation in point slope form: \[y - y_0 = m(x-x_0)\]